Please help me solve this problem: The measure of an angle is 4 less than 5 times the other. The sum of the angles is 124.
Answer by macston(5194)

.
x=second angle; 5x-4=first angle
.
(5x-4)+(x)=124
6x=128
x=21.33 The measure of the second angle is 21.33.
5x-4=5(21.33)-4=106.67-4=102.67
The first angle is 102.67.
.
CHECK:
Angle 1 + Angle 2 = 124
102.67 + 21.33 = 124
124 = 124